OpenStudy (happykiddo):

A polyhedron has 7 faces and 12 edges. What is the number of vertices of the polyhedron?

OpenStudy (happykiddo):

help

OpenStudy (anonymous):

V + F -E = 2 V + 7 - 12 = 2 V = 2+ 5 V= 7

OpenStudy (anonymous):

I used Euler's Formula.

OpenStudy (happykiddo):

very nice yes thank you

OpenStudy (anonymous):

You're welcome! I just re-read my replies and I noticed I sound so unfriendly! Sorry about that. Have a great day :)